# Do LLMs estimate uncertainty well in instruction-following?

### v3

@@ Abstract
Large language models (LLMs) could be valuable personal AI agents across various domains, provided they can precisely follow user instructions. However, recent studies have shown significant limitations in LLMs’ instruction-following capabilities, raising concerns about their reliability in high-stakes applications. Accurately estimating LLMs’ uncertainty in adhering to instructions is critical to mitigating deployment risks. We present, to our knowledge, the first systematic evaluation of uncertainty estimation abilities of LLMs in the context of instruction-following. Our study identifies key challenges with existing instruction-following benchmarks, where multiple factors are entangled with uncertainty stemming from instruction-following, complicating the isolation and comparison across methods and models. To address these issues, we introduce a controlled evaluation setup with two benchmark versions of data, enabling comprehensive comparison of uncertainty estimation methods under various conditions. Our findings show that existing uncertainty methods struggle, particularly when models make subtle errors in instruction following. While internal model states provide some improvement, they remain inadequate in more complex scenarios. The insights from our controlled evaluation setups provide crucial understanding of LLMs’ limitations and potential for uncertainty estimation in instruction-following tasks, paving the way for more trustworthy AI agents. [^1] We will release the data and code on GitHub upon publication.
@@ 2.1 Methods
Baseline uncertainty estimation methods To evaluate uncertainty in instruction-following, we employ several baseline methods, including self-evaluation of their own uncertainty (verbalized confidence, normalized p(true) and p(true)) and logits-based method (perplexity, sequence probability, and mean token entropy):
@@ 2.3 Challenges in evaluating uncertainty estimation using existing datasets
Figure 2: Existing instruction-following datasets only evaluate uncertainty estimation methods in length-biased settings, missing comparisons on controlled, length-neutral conditions. (a) Token lengths distribution for LLaMA-2-chat-7B shows that incorrect responses tend to be longer than correct ones. This length signal is prevalent in existing dataset like IFEval, where naturally generated responses are used. (b) Token length differences broken down by instruction type and model, with positive values showing that incorrect responses tend to be longer. However, this trend is not consistent across all instructions or models, highlighting the need for controlled evaluation setups. (c) Token length distribution in the Controlled version of our benchmark, where token length is balanced across correct, incorrect, and subtly incorrect responses. This setup allows us to evaluate uncertainty estimation in both length-biased and length-neutral settings.
@@ 3.2 Results on our benchmark data
3.2.1 Comparison of uncertainty methods
Controlled-Easy: Self-evaluation methods outperform logit-based ones In simpler tasks, self-evaluation methods such as verbalized confidence (short answer) and normalized-p(true) (binary choice) consistently outperformed logit-based approaches like perplexity, sequence probability, and entropy. Furthermore, for models like LLaMA2-chat-7B, LLaMA2-chat-13B, and Mistral-7B-Instruct, verbalized confidence performed better than normalized p(true), meaning short-form answer verbalized scores were more calibrated than binary choices in these cases. This overall trend indicates that for Controlled-Easy tasks, models are better at assessing their own correctness using self-evaluation methods compared to logit-based uncertainty estimation.
Controlled-Hard: Mixed performance between logit-based methods and self-evaluation methods In the instruction-following samples with more nuanced mistakes, there was a mixed performance between logit-based methods and Normalized p(true). For LLaMA2-chat-7B and LLaMA2-chat-13B, logit-based methods like perplexity and entropy outperformed verbalized confidence and normalized p(true), suggesting that as task complexity increased, logit-based methods became more reliable with those models. However, for Mistral-7B-Instruct and Phi-3-mini, normalized p(true) continued to provide better uncertainty estimation than logit-based methods, demonstrating that the best method can vary depending on the model and its underlying architecture in Controlled-Hard tasks. Uncertainty estimation scores in Controlled-Hard were consistently lower than in other set- tings and no methods had reliably estimated uncertainty in this setting.
Realistic: Sequence probability excels, but normalized p(true) remains a strong contender In the Realistic evaluation, where models were evaluated on responses with realistic patterns (including length effect), sequence probability performed best across all models, likely due to its ability to exploit the length effect in incorrect responses. However, aside from sequence probability, normalized p(true) consistently ranked as the second-best method across all models, outperforming other logit-based methods like perplexity and mean token entropy. This demonstrates that while sequence probability can take advantage of length bias, normalized p(true) offers a more balanced and reliable uncertainty estimation when length effect is less prominent.
@@ 3.2.3 Comparison of Instruction Types
4 Do internal states of LLMs know more about their uncertainty?
In the previous section, we demonstrated that both self-evaluation and logit-based methods performed poorly in the Controlled-Hard setting, where instruction-following mistakes were subtle. AUROC for uncertainty estimation was often below 0.6. Drawing inspiration from Liu et al. (2024), which suggests that model representations contain valuable information for uncertainty estimation in tasks like question and answering, we investigated whether this internal knowledge could similarly improve uncertainty estimation in instruction-following tasks. We trained a linear model as an uncertainty estimation function that maps the internal representations of LLMs to instruction-following success labels, where the probability predicted by the linear model used as uncertainty scores. The ground truth for success or failure is determined by a deterministic program that verifies instruction adherence using the IFEval dataset. We refer to this as Probe.
4.1 Settings
Representations We analyzed four language models: LLaMA-2-7B-chat, LLaMA-2-13B-chat, Mistral-7B-Instruct-v0.3, and Phi-3-mini-128k-instruct. For each model, we collected representations across three different layers: early, middle, and the final layer. Specifically, we use layers 16, 32, and 40 and for LLaMA-2-13B-chat and 14, 26, and 32 for other three models.
Linear Model and metric We train a linear model on representations on instruction-following success labels, optimized with AdamW, a 0.001 learning rate, 0.1 weight decay. The model is trained for 1000 epochs on 70% training set and is evaluated on 30% test set. We use AUROC as metric.
4.2 Results
Middle layers consistently offer the most informative representations for uncertainty estimation, particularly in simpler tasks (Controlled-Easy). Appendix Table 9 shows how Probe performance varies across instruction types and layers within the LLMs. The middle layers consistently provide the best signals for uncertainty, indicating that future work could focus on extracting more refined signals from these middle layers to enhance uncertainty estimation methods.
@@ A.3 Prompts for generating our Controlled benchmark dataset
A.4 Stats of our benchmark data
@@
Figure 3: Token length distributions for the Controlled and Realistic versions of our benchmark dataset. (a) Token length distribution in the Controlled version, where token lengths are carefully balanced between correct and incorrect responses. (b) Token length distribution in the Realistic version, where token length reflects the natural variability of model-generated responses.
@@
A.5 More results on our benchmark data
@@
A.6 More results on internal states of LLMs
@@
A.7 Length effect models
@@
Figure 10: Token length distributions for LLaMA-2-chat-13B, Mistral-7B-Inst-v0.3, and Phi-3-128k-inst. The distributions show that incorrect responses tend to be longer than correct ones, a pattern observed in existing datasets like IFEval, where naturally generated responses are used. This length signal highlights the potential bias in uncertainty estimation methods that are sensitive to token length.

### v4

@@ Abstract
Large language models (LLMs) could be valuable personal AI agents across various domains, provided they can precisely follow user instructions. However, recent studies have shown significant limitations in LLMs’ instruction-following capabilities, raising concerns about their reliability in high-stakes applications. Accurately estimating LLMs’ uncertainty in adhering to instructions is critical to mitigating deployment risks. We present, to our knowledge, the first systematic evaluation of uncertainty estimation abilities of LLMs in the context of instruction-following. Our study identifies key challenges with existing instruction-following benchmarks, where multiple factors are entangled with uncertainty stemming from instruction-following, complicating the isolation and comparison across methods and models. To address these issues, we introduce a controlled evaluation setup with two benchmark versions of data, enabling comprehensive comparison of uncertainty estimation methods under various conditions. Our findings show that existing uncertainty methods struggle, particularly when models make subtle errors in instruction following. While internal model states provide some improvement, they remain inadequate in more complex scenarios. The insights from our controlled evaluation setups provide crucial understanding of LLMs’ limitations and potential for uncertainty estimation in instruction-following tasks, paving the way for more trustworthy AI agents. [^1] Code and data are available at https://github.com/apple/ml-uncertainty-llms-instruction-following
Work done while at Apple.
@@ 2.1 Methods
Baseline uncertainty estimation methods To evaluate uncertainty in instruction-following, we employ several baseline methods, including self-evaluation of their own uncertainty (verbalized confidence, normalized p(true) and p(true)), logits-based method (perplexity, sequence probability, and mean token entropy), and probing method:
@@
  • Probing: Drawing inspiration from Liu et al. (2024), which suggests that model representations contain valuable information for uncertainty estimation in tasks like question and answering, we investigated whether this internal knowledge could similarly improve uncertainty estimation in instruction-following tasks. We trained a linear model as an uncertainty estimation function that maps the internal representations of LLMs to instruction-following success labels, where the probability predicted by the linear model used as uncertainty scores. We use layers 16, 32, and 40 and for LLaMA-2-13B-chat and 14, 26, and 32 for other three models. The ground truth for success or failure is determined by a deterministic program that verifies instruction adherence using the IFEval dataset. We train a linear model on representations on instruction-following success labels, optimized with AdamW, a 0.001 learning rate, 0.1 weight decay. The model is trained for 1000 epochs on 70% training set and is evaluated on 30% test set. We refer to this as Probe.
@@ 2.3 Challenges in evaluating uncertainty estimation using existing datasets
Figure 2: Existing instruction-following datasets only evaluate uncertainty estimation methods in length-biased settings, missing comparisons on controlled, length-neutral conditions. The distributions are normalized by the total number of responses in each class. (a) Token lengths distribution for LLaMA-2-chat-7B shows that incorrect responses tend to be longer than correct ones. This length signal is prevalent in existing dataset like IFEval, where naturally generated responses are used. (b) Token length differences broken down by instruction type and model, with positive values showing that incorrect responses tend to be longer. However, this trend is not consistent across all instructions or models, highlighting the need for controlled evaluation setups. (c) Token length distribution in the Controlled version of our benchmark, where token length is balanced across correct, incorrect, and subtly incorrect responses. This setup allows us to evaluate uncertainty estimation in both length-biased and length-neutral settings.
@@ 3.2 Results on our benchmark data
3.2.1 Comparison of uncertainty methods: Controlled-Easy and Realistic
In Probe, middle layers consistently offer the most informative representations for uncertainty estimation, particularly in simpler tasks (Controlled-Easy). Appendix Table 10 shows how Probe performance varies across instruction types and layers within the LLMs. The middle layers consistently provide the best signals for uncertainty, indicating that future work could focus on extracting more refined signals from these middle layers to enhance uncertainty estimation methods.
Self-evaluation methods outperform logit-based ones in Controlled-Easy In simpler tasks, self-evaluation methods such as verbalized confidence (short answer) and normalized-p(true) (binary choice) consistently outperformed logit-based approaches like perplexity, sequence probability, and entropy. Furthermore, for models like LLaMA2-chat-7B, LLaMA2-chat-13B, and Mistral-7B-Instruct, verbalized confidence performed better than normalized p(true), meaning short-form answer verbalized scores were more calibrated than binary choices in these cases. This overall trend indicates that for Controlled-Easy tasks, models are better at assessing their own correctness using self-evaluation methods compared to logit-based uncertainty estimation.
Sequence probability excels, but normalized p(true) remains a strong contender in Realistic In the Realistic evaluation, where models were evaluated on responses with realistic patterns (including length effect), sequence probability performed best across all models, likely due to its ability to exploit the length effect in incorrect responses. However, aside from sequence probability, normalized p(true) consistently ranked as the second-best method across all models, outperforming other logit-based methods like perplexity and mean token entropy. This demonstrates that while sequence probability can take advantage of length bias, normalized p(true) offers a more balanced and reliable uncertainty estimation when length effect is less prominent.
3.2.2 Comparison of uncertainty methods: Controlled-Hard
Mixed performance between logit-based methods and self-evaluation methods in Controlled-Hard In the instruction-following samples with more nuanced mistakes, there was a mixed performance between logit-based methods and Normalized p(true). For LLaMA2-chat-7B and LLaMA2-chat-13B, logit-based methods like perplexity and entropy outperformed verbalized confidence and normalized p(true), suggesting that as task complexity increased, logit-based methods became more reliable with those models. However, for Mistral-7B-Instruct and Phi-3-mini, normalized p(true) continued to provide better uncertainty estimation than logit-based methods, demonstrating that the best method can vary depending on the model and its underlying architecture in Controlled-Hard tasks. Uncertainty estimation scores in Controlled-Hard were consistently lower than in other set- tings and no methods had reliably estimated uncertainty in this setting.
@@ 5 Conclusion
Acknowledgments
This work was conducted during an internship at Apple AIML. We sincerely thank Sinead Williamson and Udhay Nallasamy for their valuable feedback and insightful suggestions on this work. We are also grateful to Guillermo Sapiro for his unwavering support and guidance throughout the research.
@@ A.3 Prompts for generating our Controlled benchmark dataset
Prompt for generating subtle off-target responses {Instruction-and-task input}
@@
A.4 Why do we choose IFEval dataset?
Here, we would like to emphasize why we choose IFEval as our primary dataset instead of using real-world dataset with different contexts and domains.
First, we select IFEval to focus on our scope which is ‘single, simple, and non-ambiguous instructions’. Real-world datasets often involve complex, ambiguous, or multi-instruction prompts, which can conflate multiple factors affecting uncertainty. For this first systematic evaluation, we chose to focus on single, simple, and verifiable instructions to ensure clarity and isolate the uncertainty estimation process. The IFEval dataset is well-suited for this purpose, as it provides 25 distinct types of simple and clear instructions that align with our goal of establishing a robust baseline.
Second, we want to avoid evaluator-induced uncertainties. Most real-world tasks and benchmark datasets rely on LLM-based evaluators to determine whether a response follows an instruction. However, LLM-based evaluators may introduce their own uncertainties or make errors in assessing success or failure, which could obscure the true uncertainty estimation capabilities of the tested models. The IFEval dataset avoids this issue by including instructions with deterministic evaluation programs that objectively verify compliance. For instance, an instruction like “please do not include keywords: …” can be automatically validated using a simple program to check for the presence of those keywords. This feature eliminates ambiguity in evaluation and allows us to directly focus on LLMs’ ability to estimate uncertainty.
Our main contribution is the careful design of benchmark data specifically tailored to ‘uncertainty estimation’ in instruction-following contexts. We believe that those underlying considerations (e.g., length bias, task quality entanglement, and varying levels of model difficulty) presented in our study can extend to future instruction-following datasets. While IFEval serves as an ideal starting point for this research, we hope our framework inspires future efforts to tackle uncertainty estimation in more complex, real-world tasks.
A.5 IFEval data examples
Table 6 presents examples from the IFEval dataset, such as tasks like writing a resume or creating a joke about programmers. The instructions assigned to each task vary, requiring the model to follow specific guidelines such as including or excluding certain keywords, ensuring word usage meets a specific frequency, and adhering to formatting rules. The keywords that must be included or excluded differ based on the task. For instance, in the resume task, keywords might include “resume”, “software”, or “engineer”, whereas in the joke task, the focus may shift to terms like “syntax” or “code”. These varied instructions introduce diverse challenges for the model in instruction-following.
Type Example
Task Write a resume for a software engineer with 5+ years of experience in the Bay Area, CA.
Instruction
keywords:existence
Make sure to include the keywords: “skills”, “technology”, “career”.
keywords:forbidden
Do not include the following keywords: resume, software, engineer, experience.
keywords:frequency
Make sure to use the word “qualifications” at least 2 times.
startend:end checker
Your resume must end with the exact phrase “Looking forward to contributing to innovative projects.”
detectable content:number placeholders
Make sure to include at least 5 placeholders represented by square brackets, such as [name].
Task Write a joke about programmers.
Instruction
keywords:existence
Make sure to include the keywords: “humor”, “code”, “life”.
keywords:forbidden
Do not include the following keywords: joke, programmers.
keywords:frequency
Make sure to use the word “syntax” at least 3 times.
startend:end checker
Your programmer joke must end with the exact phrase “And that’s the real bug in the code of life.”
detectable content:number placeholders
Make sure to include at least 3 placeholders represented by square brackets, such as [name].
Table 6: Examples from the IFEval dataset. This table shows two tasks: writing a resume and crafting a joke about programmers. Each task is paired with multiple instruction types, such as including/excluding keywords, ensuring word frequency, and adhering to specific content formatting rules.
A.6 Stats of our benchmark data
@@
Figure 3: Token length distributions for the Controlled and Realistic versions of our benchmark dataset. The distributions are normalized by the total number of responses in each class. (a) Token length distribution in the Controlled version, where token lengths are carefully balanced between correct and incorrect responses. (b) Token length distribution in the Realistic version, where token length reflects the natural variability of model-generated responses.
@@
A.7 More results on our benchmark data
@@
A.8 More results on internal states of LLMs
@@
A.9 Length effect models
@@
Figure 10: Token length distributions for LLaMA-2-chat-13B, Mistral-7B-Inst-v0.3, and Phi-3-128k-inst. The distributions are normalized by the total number of responses in each class. The distributions show that incorrect responses tend to be longer than correct ones, a pattern observed in existing datasets like IFEval, where naturally generated responses are used. This length signal highlights the potential bias in uncertainty estimation methods that are sensitive to token length.